# Nemotron-CrossThink: Scaling Self-Learning beyond Math Reasoning

### v2

@@ Appendix B Token Efficiency Analysis
Task Type Model Min Max Mean
GPR $\mathcal{B}_{gpr\uparrow}$ 83.20 2697.80 385.41
$\mathcal{B}_{only\_math}$ 159.60 9594.00 638.57
orz 223.00 8221.80 1114.60
Math $\mathcal{B}_{gpr\uparrow}$ 170.25 10130.00 622.00
$\mathcal{B}_{only\_math}$ 201.75 11330.25 730.68
orz 292.00 12917.00 1257.00
Across gpr tasks, $\mathcal{B}_{gpr\uparrow}$ produces the most concise correct responses, with a mean of 385 tokens—39.6% fewer than $\mathcal{B}_{only\_math}$ and 65.4% fewer than orz. This suggests that training with multi-domain corpora equips the model to reason more efficiently in less structured tasks, avoiding unnecessarily verbose responses.
@@
Figure 4: Average token lengths of correct and incorrect responses across general-purpose and math reasoning tasks for models trained on $\mathcal{B}_{gpr\uparrow}$, $\mathcal{B}_{only\_math}$, and orz.
@@
Recent studies such as DeepScaler (Luo et al., 2025) have noted that incorrect answers often exhibit longer trajectories, leading to wasted computation and less efficient learning. Echoing this observation, we analyze the average token lengths of correct and incorrect responses for models trained on different blends: $\mathcal{B}_{gpr\uparrow}$, $\mathcal{B}_{only\_math}$, and orz.
As shown in Figure 4, incorrect responses are consistently and substantially longer than correct ones—by 3.6$\times$ on average. This pattern holds across both general-purpose and math reasoning tasks, suggesting that verbose reasoning does not guarantee correctness. In fact, longer responses often reflect the model’s uncertainty, overthinking, or repetitive CoT traces, rather than productive deduction.
Appendix C Sub-category Accuracy Analysis
@@
Figure 5: Sub-category Accuracy Comparison across mmlu-pro Domains. The $\mathcal{B}_{gpr\uparrow}$ blend consistently outperforms $\mathcal{B}_{only\_math}$ in a wide range of non-math reasoning categories such as business, law, psychology, and economics. Surprisingly, it also slightly surpasses the math-specialized blend in the mmlu-pro math category, highlighting the generalizability and versatility of multi-domain training.
Figure 6: Sub-category Accuracy Comparison across agieval. While $\mathcal{B}_{only\_math}$ performs marginally better in the math, $\mathcal{B}_{gpr\uparrow}$ achieves stronger results in non-math domains.
@@
Figure 7: Sub-category Accuracy Comparison across supergpqa. The $\mathcal{B}_{gpr\uparrow}$ blend consistently outperforms $\mathcal{B}_{only\_math}$ in a wide range of non-math reasoning categories except the science category which consists of fields like mathematics, physics, astronomy, chemistry etc.—highlighting the generalizability and versatility of multi-domain training.

### v3

@@ Appendix E Token Efficiency Analysis
Across gpr tasks, $\mathcal{B}_{gpr\uparrow}$ produces the most concise correct responses, with a mean of 385 tokens—39.6% fewer than $\mathcal{B}_{only\_mr}$ and 65.4% fewer than orz. This suggests that training with multi-domain corpora equips the model to reason more efficiently in less structured tasks, avoiding unnecessarily verbose responses.
@@
Figure 7: Average token lengths of correct and incorrect responses across general-purpose and math reasoning tasks for models trained on $\mathcal{B}_{gpr\uparrow}$, $\mathcal{B}_{only\_mr}$, and orz.
@@
Recent studies such as DeepScaler (Luo et al., 2025) have noted that incorrect answers often exhibit longer trajectories, leading to wasted computation and less efficient learning. Echoing this observation, we analyze the average token lengths of correct and incorrect responses for models trained on different blends: $\mathcal{B}_{gpr\uparrow}$, $\mathcal{B}_{only\_mr}$, and orz.
[…]
Appendix F Sub-category Accuracy Analysis
@@
Figure 8: Sub-category Accuracy Comparison across mmlu-pro Domains. The $\mathcal{B}_{gpr\uparrow}$ blend consistently outperforms $\mathcal{B}_{only\_mr}$ in a wide range of non-math reasoning categories such as business, law, psychology, and economics. Surprisingly, it also slightly surpasses the math-specialized blend in the mmlu-pro math category, highlighting the generalizability and versatility of multi-domain training.
Figure 9: Sub-category Accuracy Comparison across agieval. While $\mathcal{B}_{only\_mr}$ performs marginally better in the math, $\mathcal{B}_{gpr\uparrow}$ achieves stronger results in non-math domains.
@@
Figure 10: Sub-category Accuracy Comparison across supergpqa. The $\mathcal{B}_{gpr\uparrow}$ blend consistently outperforms $\mathcal{B}_{only\_mr}$ in a wide range of non-math reasoning categories except the science category which consists of fields like mathematics, physics, astronomy, chemistry etc.—highlighting the generalizability and versatility of multi-domain training.
@@
Appendix G Generalizability Across Model Architectures
To verify that the efficacy of Nemotron-CrossThink is not limited to a single model architecture, we extended our evaluation to the Nemotron-H (NVIDIA et al., 2025) model family. We conducted a controlled experiments only for 200 steps using our two best performing data blends: $\mathcal{B}_{only\_mr}$, a domain-specific “Math-Only” blend and $\mathcal{B}_{gpr\uparrow}$, the “CrossThink” mixed-domain blend (utilizing a 2:1 ratio of general-purpose to math domains).
$\mathcal{B}_{only\_mr}$ 68.48 48.58 34.01 57.94 23.10 75.36 51.50 46.42
$\mathcal{B}_{gpr\uparrow}$ 69.94 49.66 39.09 58.23 24.60 70.88 48.17 48.30
Table 13: Expanding to $\mathcal{M}$=Nemotron-H: The relative trend confirms that Nemotron-CrossThink improves general reasoning (MMLU, GPQA, etc.) compared to the Math-Only baseline.
As observed in Table 13, the CrossThink blend demonstrates a strong capacity for transfer learning. While the math-specific scores (math-500, amc23) are slightly lower than the Math-Only baseline—an expected outcome given that the CrossThink blend contains fewer math samples due to the 2:1 mixing ratio—the model significantly gains in general reasoning capabilities. Specifically, CrossThink outperforms the baseline on general-purpose reasoning tasks, achieving a 1.88% improvement in the Reasoning Average. This suggests that mixing domains helps the model generalize better to unseen reasoning tasks, even within a new architecture.
Beyond performance metrics, we analyzed the inference efficiency of the approach on the Nemotron-H architecture. We calculated the average number of tokens generated for correct responses across seven benchmarks.
Task $\mathcal{B}_{only\_mr}$ $\mathcal{B}_{\uparrow gpr}$ Reduction
mmlu 387.69 313.39 19.2%
agieval 525.10 432.25 17.7%
mmlu-pro 772.27 495.56 35.8%
gpqa-diamond 1124.62 656.15 41.7%
supergpqa 421.52 306.13 27.4%
math-500 551.96 460.72 16.5%
amc23 557.42 438.66 21.3%
Average 620.08 443.26 28.5%
Table 14: Average token counts for correct responses (Nemotron-H). Nemotron-CrossThink significantly reduces the verbosity of the model across all benchmarks, achieving a 28.5% average reduction in token generation for correct answers.
As shown in Table 14, the Nemotron-CrossThink model generates 28.51% fewer tokens on average compared to the baseline while maintaining or improving accuracy on general reasoning tasks. This confirms that the efficiency gains observed in Qwen-2-7B are architectural-agnostic: Nemotron-CrossThink consistently encourages more concise reasoning paths without sacrificing solution correctness.
Appendix H Additional Discussion
Why Nemotron-CrossThink is applied directly on the base model?
In this work, we apply RL directly to pretrained language models rather than to fine-tuned models. Although RL on SFT models is often considered more stable due to instruction-following priors and task structure acquired during supervised training (DeepSeek-AI, 2025), prior findings show that this choice does not lead to consistent benefits.
Several studies report that SFT can reduce model entropy, which is important for effective exploration during reinforcement learning (Chu et al., 2025), and that applying RL methods such as GRPO (Shao et al., 2024) to already aligned or instruction-tuned models can hurt performance (Chen et al., 2025). Other work observes improvements when RL is applied after heavy SFT (Liu et al., 2025). These mixed outcomes indicate high variance in the effectiveness of RL on SFT models, making it difficult to isolate the contribution of reinforcement learning itself.
We therefore focus on RL applied to the pretrained base model to study the emergence of reasoning abilities without strong instruction-following priors and to avoid confounding effects introduced by SFT. This choice is further motivated by recent work showing that self-learning on base models can yield substantial gains in reasoning capabilities across diverse tasks (Zeng et al., 2025b; Hu et al., 2025b; Wang et al., 2025).
Robustness across Decoding Strategies (Greedy vs. Pass@1[8]).
In addition to greedy decoding, we evaluated two of our strongest models using pass@1[8] with stochastic decoding (temperature = 0.6, top_p = 0.95, k = 8). The results are summarized in Table 15.
Model mmlu mmlu-pro gpqa-diamond agieval supergpqa math-500 amc23 aime24 Avg
$\mathcal{B}_{only\_mr}$ 72.26 53.40 34.54 61.62 22.35 76.88 62.00 16.68 49.97
$\mathcal{B}_{gpr\uparrow}$ 73.82 57.01 37.86 64.00 25.14 76.03 61.11 15.27 51.28
Table 15: Pass@1[8] performance with stochastic decoding (temperature = 0.6, top_p = 0.95, $k=8$).
Under pass@1[8], the $\mathcal{B}_{gpr\uparrow}$ model achieves the highest overall average score, outperforming the $\mathcal{B}_{only\_mr}$ model. The gains are most pronounced on general-purpose reasoning benchmarks, including mmlu, mmlu-pro, gpqa-diamond, etc., where $\mathcal{B}_{gpr\uparrow}$ consistently improves over $\mathcal{B}_{only\_mr}$. On math-focused tasks such as math-500, amc23, and aime24, the two models exhibit comparable performance, with $\mathcal{B}_{only\_mr}$ showing a small advantage on some benchmarks. These results mirror the trends observed with greedy decoding and further support that the multi-domain blend improves broad reasoning performance without sacrificing math capability.
Appendix I Extended Related Work
Self-Learning beyond Math.
High-quality training data are crucial for scalable Reasoner-Zero training. Most of the recent works emphasize mathematical benchmark-centric data (AMC, AIME, Math, Olympiads, and AoPS) for reinforcement learning (Hu et al., 2025a; Aggarwal and Welleck, 2025; Trung et al., 2024; Ye et al., 2025; Zeng et al., 2025a) as designing verifiable rewards is much easier for math tasks. They exclude problems such as multiple choice and proof-oriented problems which reduces the answer space diversity. mcq type of questions are important for mmlu and other non-reasoning centric tasks. Recently, Ma et al. (2025); Su et al. (2025) attempted to address this with a model-based verifier to handle diversity in the answer space. However, as discussed in previous works, llm-as-a-Judge may suffer from pitfalls of reward hacking (DeepSeek-AI, 2025; Weng, 2024; Wen et al., 2024) and further diverge the model more from the correct reasoning processes. Additionally, they do not show any analysis to estimate the contribution of each domain in the final task performance. Despite training on diverse domains, Nemotron-CrossThink offers simple, scalable and robust reward estimation without any external reward model. For a rule-based reward model, the format of input data and the final answer is crucial and largely underexplored. Furthermore, their additional sources of data synthesis approach has no details making it infeasible to scale for domains other than math. The kind of data and the ratio of each type of data important for the overall improvement of llms across multiple benchmarks have yet to be explored.